OpenStudy (anonymous):

A cannery processed 605 pounds of strawberries in 3.5 hours. The cannery processed 2100 pounds in 10 hours. How many pounds of strawberries can be processed in 12 hours?

OpenStudy (anonymous):

2100 lbs/10hours = 210lb/1hours = 2720 lbs/12hours

OpenStudy (tkhunny):

Let's see... Model: Pounds = k*Hours 605/3.5 = 172.85714285714285714285714285714 2100/10 = 210 Okay, the simplest linear model will not do. Model: Pounds = A + k*Hours 605 = A + k(3.5) 2100 = A + k(10) 1495 = k(6.5) ==> k = 230 <== Looks promising. 2100 = A + 230(10) ==> A = -200 <== Well, that doesn't make any sense!! They produce -200 every zero (0) hours? Model: Pounds = Ae^(kx) <== A simple exponential model. 605 = Ae^(k3.5) 2100 = Ae^(k10) 2100/605 = e^(10k)/e^(3.5k) 3.47107438 = e^(6.5k) ==> k = 0.1914560 605 = Ae^(0.1914560(3.5)) ==> A = 309.55397 That looks okay. 309.55397e^(0.1914560(12)) = 3,080 or so. That model may have some application to the problem. Allofus seems to be ignoring half the data. Don't do that. Find a model that seems reasonable and go with it.